OpenStudy (anonymous):

If one micrometeorite (a sphere with a diameter of 1.4 10-6 m) struck each square meter of the Moon each second, estimate the number of years it would take to cover the Moon with micrometeorites to a depth of one meter? Hint: Consider the volume of a micrometeorite and a cubic box, 1 m on a side, on the Moon. Find how long it would take to fill the box. (Type your answer using one of the following formats, 1.2e-3 for 0.0012 and 1.20e+2 for 120)

OpenStudy (anonymous):

\[Moon surface = 3.793 × 10^7 km^2 = 3.793 × 10^{13} m^2=boxes\] The volume of 1 box = 1 m^3 The volume of 1 sphere = \[4/3 \pi r ^{3} = 4/3 \pi 0.7 × 10^{-6} = 2.9 × 10^{-6} m^3\] So 1 box holds \[ 1 / ( 2.9 * 10^{-6}) = 344828\] Each box is 'filled' with 1 meteorite every second, so it'd take 344828 seconds = 0.01093 years = \[1.093 × 10^{-2} years\]